Pinnocle Corporation has provided the following data from its activity-based costing system:

Activity Cost Pool Total Cost Total Activity
Assembly $ 1,384,600 70,000 machine-hours
Processing orders $ 48,006 1,800 orders
Inspection $ 127,917 1,830 inspection-hours

The company makes 500 units of product S78N a year, requiring a total of 750 machine-hours, 40 orders, and 10 inspection-hours per year. The product's direct materials cost is $50.69 per unit and its direct labor cost is $12.78 per unit. The product sells for $116.20 per unit.

According to the activity-based costing system, the product margin for product S78N is:

Multiple Choice

  • $26,365.00.

  • $10,831.00.

  • $10,463.20.

  • $9,764.20.
Answer #1

Direct material cost = 500 X $50.69 = $25,345

Direct labor cost = 500 X $12.78 = $6,390

Assembly cost = ($1,384,600 / 70,000) X 750 = $14,835

Processing order costs = ($48,006 / 1,800) X 40 = $1,066.8

Inspection costs = ($127,917 / 1,830) X 10 = $699

Total cost = $48,335.8

Total sales = 500 X $116.20 = $58,100

Product margin = $9,764.2

4th option